OpenStudy (anonymous):

What is the slope of a line parallel to the line with equation 5x + 3y =7? A. 5/3 B. -7/3 C. -5/3 D. -3/5

OpenStudy (anonymous):

@nirmalnema

OpenStudy (leader):

3y=-5x+7 y=-5/3+7/3

DivineSolar (divinesolar):

First of all lets walk you through this. What do we do first to solve the equation

OpenStudy (leader):

So your answer is C -5/3

DivineSolar (divinesolar):

Leader. Read the moderators comment :/

OpenStudy (nirmalnema):

-5/3

DivineSolar (divinesolar):

You guys are breaking openstudys polices.

DivineSolar (divinesolar):

policies.

OpenStudy (anonymous):

thanks everyone @=

DivineSolar (divinesolar):

Let's solve for x. 5x+3y=7 Step 1: Add -3y to both sides. 5x+3y+−3y=7+−3y 5x=−3y+7 Step 2: Divide both sides by 5. 5x 5 = −3y+7 5 x= −3 5 y+ 7 5

OpenStudy (nirmalnema):

slope=-a/b (ax+by+c=0))

DivineSolar (divinesolar):

so -3/5